OpenStudy (anonymous):

Will fan and medal please help!!! What is the formula for calculating the length of the space diagonal of a right rectangular prism on the basis of the prism's length L, width W, and height H? A. d = L² + w² + h² B. h² = L² + w² C. d = √L + W + H D. d = √L² + W² + H² What is the volume of a prism that has a base area of 24 cm² and a height of 12 cm? A. 336 cm³ B. 576 cm³ C. 288 cm³ D. 216 cm³ What are the coordinates of point R? A. (0, -2, 5) B. (2, -5, 0) C. (0, 2, 5) D. (-2, 5, 0) What is the distance between (5, 6, 4) and (4, -4, -2)? Round to the nearest tenth of a uni

candycove (candycove):

:')

OpenStudy (anonymous):

Hi Candy. Can you help me still? https://learning.k12.com/d2l/le/content/109854/viewContent/17304849/View

imqwerty (imqwerty):

1st one-  1st you see that triangle ABC its a right triangle and we can find the diagonal AC by pythagorus theorem it will be \(AC^2=AB^2+BC^2\) \(AC^2=l^2+w^2\) \(AC= \sqrt{l^2+w^2}\) now you see that triangle ACD apply pythagorus theorem again \(AD^2=AC^2+CD^2\) AD is the diagonal which we need \(AD^2=l^2+w^2+h^2\)